If log 8 base x +   log  base 8 = 10/3, find  value 
Of x

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Let assume
Log x 8 = t
 
Now
 
t + 1/t = 10/3
 
3t² - 10t + 3 = 0
 
t = 3
Then
log x 8 = 3
 
x = 2
 
 
Now if t = 1/3
 
log 8 = 1/3
 
x = 512
 
 
Regards
Arun